OpenStudy (anonymous):

Please need an explanation!:)))) I WILL MEDAL 2/1= 8X-2/9

OpenStudy (campbell_st):

so is the question \[\frac{2}{1} = \frac{8x - 2}{9}\]

OpenStudy (campbell_st):

or \[\frac{2}{1} = 8x - \frac{2}{9}\]

OpenStudy (anonymous):

This link tells you everything about the question :) @beauty_queen67 http://www.wolframalpha.com/input/?i=2%2f1%3d+8x-2%2f9&incTime=true